solve the equation show as decimal 0.0000
4(3^x) = 5

Answer by Alan3354(69443)

4(3^x) = 5
3^x = 5/4
log(3^x) = log(1.25)
3log(x) = log(1.25)
log(x) = log(1.25)/3 =~ 0.03230337
x = 10^0.323...
x =~ 1.077217